# MS-IMAP - A Multi-Scale Graph Embedding Approach for Interpretable Manifold Learning

### v4

@@ 3.1 Step 1: Feature Representation to encode multi-scale structure
3D-Tensor Encoding: The 3D-Tensor Encoding method concatenates the multi-scale representation for all features at a fixed scale, generating a matrix representation $\psi_{\mathbf{x}}(s_{j},:,:)\in\mathbb{R}^{D}\times\mathbb{R}^{N}$ for each scale $s_{j}$. After concatenating all $f_{i}$ at a fixed scale $s_{j}$, we have
@@
[…]
$\mathbf{c}(\psi_{{f}_{i}}(s_{k},:,:),\psi_{{f}_{j}}(s_{k},:,:))=\psi_{{f}_{i}}(s_{k},:,:)\left|\right|\psi_{{f}_{j}}(s_{k},:,:)$
The optimization can be performed separately for each scale $s_{j}$, or jointly for all scales $s_{j}$ using the 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$ We detail the proposed feature representation encoding methods in the pseudo code algorithm 1.
@@
  Output: Initial Node Embeddings ${\psi}_{\mathbf{x}}$
@@
1  Step 4: Initial Embedding construction:
@@
  Step 5: Concatenate $\psi_{{f}_{r}}(s_{j},:)$ for $r=1,...,D$ in a matrix form $\psi_{\mathbf{x}}(s_{j},:,:)$ for each fixed scale $s_{j}$, representing all scales using 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
@@
3D-Tensor Based Optimization for graph Embedding
The encoded manifold representation is a 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$. An SGD update rule at iteration $t$, applied to the 3D tensor, is:
@@
where $v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}=\frac{1}{1+||\psi_{\mathbf{x}_{i}}(s_{k},:,:)-\psi_{\mathbf{x}_{j}}(s_{k},:,:)||^{2}}$. We apply the optimization based on SGD with respect to each scale $s_{k}$, where the gradient of the loss is approximated by:
@@
The optimization process uses SGD with positive and negative sampling for all scales $s_{k}$ given a pair of positive or negative pair of nodes $i,j$.
We compute the final embedding by summing up all the optimized embeddings with respect to each scale $s_{k}$
@@
Note that the dimensionality of the final embedding
$\tilde{\psi}_{\mathbf{x}}\in\mathbb{R}^{D}\times\mathbb{R}^{N}$, maintaining the same dimensionality as the original input, and thus there exists a one-to-one correspondence between the coordinates of the input features $\mathbf{x}$ and $\tilde{\psi}_{\mathbf{x}}$. This correspondence ensures that each dimension of the feature space is directly mapped to a single dimension in the embedding space $\tilde{\psi}_{\mathbf{x}}$.
@@
Our optimization process involves using SGD with positive and negative sampling, similar to recent graph embeddings methods [33, 25]. Positive edge samples are associated with attraction (first term on the right hand side), while negative samples (second term on the right hand side) refer to a pair of nodes that are not connected in the graph, which create repulsion among dissimilar points.
The optimization can be performed jointly for all scales using the 3D tensor. Specifically, our implementation employs an optimization process that operates jointly across all scales, as described in the SGD update rule for iteration $t$ in Eq. (2). At each iteration $t$, associated with either a positive sample (an edge) or a negative sample (a pair of unconnected nodes), the update rule from Eq.(2). is applied. The derivative of the loss function, as detailed in Eq. (4), is computed for all scales $(s_{1},...s_{K})$, with the sign in Eq.(2) depending on whether the sample is positive or negative. Once the optimization concludes, the final embedding is derived from the 3D tensor by summing over the scales, as specified in Eq.(6). Encoding the representation in a 3D tensor offers two key advantages. First, is the computational efficiency: the update rule at each iteration, associated with a pair of nodes, is performed jointly and simultaneously for all scales. This avoids the need for independent optimization per scale, which would significantly increase computational complexity. The second is optimization alignment: jointly optimizing across scales ensures alignment in the optimization process for all scales with respect to a fixed pair of nodes (positive or negative sample). This approach is more effective than optimizing scales independently, as it maintains consistency and improves overall performance.
In this section, we characterize the theoretical properties of the representation power of the SGW operator by considering functions sampled from the Paley-Wiener spaces [27] on combinatorial graphs. Pesenson introduced the Paley-Wiener spaces and demonstrated that Paley-Wiener functions of low type are uniquely determined by their values on certain subsets, known as uniqueness sets $U$. We show that the SGW operator can represent functions $f$ within the Paley-Wiener space more efficiently than the graph Laplacian operator $\mathbf{L}$. This efficiency is demonstrated by showing that the SGW operator is more effective in representing functions with larger bandwidth $\omega$ in the Paley-Wiener spaces (i.e., with higher frequencies) using subsets of nodes from the uniqueness sets $U$.
To characterize the representation properties of functions defined over $PW_{\omega}(G)$, we employ the notion of the $\Lambda$-set, introduced by Pesenson which is central to our investigation. Formally, the Paley-Wiener space of $\omega$-bandlimited signals is defined as $PW_{\omega}(G)=\left\{f\mid\hat{f}(\lambda)=0\,\,\forall\,\,\lambda>\omega\right\}$. The space $L_{2}(G)$ is defined as the Hilbert space of all complex-valued functions, and $L_{2}(S)$ is defined as the space of all functions from $L_{2}(G)$ with support in $S$: $L_{2}(S)=\left\{\varphi\in L_{2}(G)\mid\varphi(v)=0,v\in V(G)\setminus S\right\}$.
The $\Lambda$-set is defined as follows: a set of vertices $S\subset V$ is a $\Lambda(S)$-set if any $\varphi\in L_{2}(S)$ satisfies the inequality $||\varphi||\leq\Lambda||\mathbf{L}\varphi||$ for some constant $\Lambda(S)>0$. The infimum of all $\Lambda>0$ for which $S$ is a $\Lambda$-set is denoted by $\Lambda$.
The ability of the SGW operator to efficiently represent functions $f\in PW_{\omega}(G)$ can be summarized in the following theorem, which highlights the role of the $\Lambda_{\psi}$-set with respect to the operator $\psi$. We show that any signal $f\in PW_{\omega}(G)$, where $\lambda_{1}\leq\omega<\Omega_{G}$ for some $\Omega_{G}<\lambda_{N}$, can be uniquely represented by its samples on the uniqueness set $U$ using the SGW operator. Under certain conditions related to the SGW operator, its associated $\Lambda_{\psi}$-set is smaller than the $\Lambda$-set associated with the Laplacian operator.
Theorem 1
Let $G=(V,\mathbf{W})$ be a connected graph with $n$ vertices and $f\in PW_{\omega}(G)$ for $\lambda_{1}\leq\omega<\lambda_{\text{max}}$. The SGW operator $\psi$ can be constructed such that the set $S$ is a $\Lambda_{\psi}$-set and the set $U=V\setminus S$ is a uniqueness set for any space $PW_{\omega}(G)$ with $\omega<1/\Lambda_{\psi}$ and $\Lambda_{\psi}<\Lambda$ for any $\varphi\in L_{2}(S)$, where $\Lambda$ is the $\Lambda$-set of the Laplacian operator.
The proof is provided in the Appendix.
5 Experimental Results
We demonstrate the advantage of our approach over existing approaches using both synthetic and real datasets. We evaluate our method by testing its clustering performance using the output embedding and comparing it to several representative methods, including UMAP, t-SNE, Diffusion Maps, ISOMAP, PHATE [26], PaCMAP [36], TriMap [1] and HeatGeo [22]. For UMAP, t-SNE, PHATE, PaCMAP, TriMap and HeatGeo, we hyperparameter tune the performance on each dataset, running the method five times for each hyperparameter configuration and taking the average. Then we report the best average score. Additional details are provided in the supplementary material.
Evaluation metrics: We test performance using the Adjusted Rand Index (ARI) and Adjusted Mutual Information (AMI). In all experiments we used $k$-means to cluster the data in the embedding space.
5.1 Real Datasets
We study the performance of MS-IMAP compared to other methods for real datasets. We chose a mix of datasets from varied fields: the Census dataset [14] is a financial dataset containing information about individuals extracted from the 1994 US Census; the Zilionis dataset [40] is a biological dataset containing single-cell sequencing data from different types of cells; the Animals with Attributes (AWA) [37] dataset is an image dataset containing images of animals; The Human Activity Recognition (HAR) [2] dataset which consists of sensor data collected from subjects performing six different activities: walking, walking upstairs, walking downstairs, sitting, standing, and lying. Each activity is represented as a feature vector of 561 attributes and 7352 samples; and the Cora dataset [30, 24] is a graph dataset of a citation network consisting of scientific publications. Note that we only use the features without the natural graph associated with Cora. More information about each dataset can be found in the supplementary material. As shown in Table 1, MS-IMAP is competitive and performs the best overall, while also providing an interpretable embedding through its one-to-one correspondence with the original features—unlike the other competing graph embedding methods.
Runtime and Computational Complexity The computational complexity of Multi-Scale IMAP is of $O(NDlog(N))$. The execution time of our method with Python code implementation with 32 cores Intel Xeon 8259CL running at 2.50Ghz and 256GB of RAM on the Cancer QC data-set of 48,969 samples and 306 features took 27 mins, on Cora 5 min, on the AWA dataset 1.35 mins and on Census 7 min. The increase in running times compared to methods such as UMAP is primarily due to our extension of manifold learning to include interpretability. Our approach generally requires the embedding dimensionality to be at least equal to the input dimensionality of the features, which adds computational complexity but ensures meaningful feature representation and interpretation. Our method is practical for handling datasets with millions of points and dozens to hundreds of features within a few hours. In fields such as finance and genomics/biology, where typical dataset sizes range from under 100,000 to several million data points, analysis is often conducted offline once the graph embedding is computed. Runtimes of up to a few hours are generally acceptable and do not present a significant challenge, making the method well-suited for real-world applications. Additional details are provided in the Appendix.
Data HAR Census Zilionis AWA Cora
Method / Accuracy ARI AMI ARI AMI ARI AMI ARI AMI ARI AMI
Features 0.45 0.58 0.11 0.15 0.48 0.63 0.07 0.16 0.01 0.07
ISOMAP 0.55 0.65 0.18 0.09 0.44 0.56 0.52 0.60 0.12 0.17
Diffusion Maps N/A N/A 0.06 0.16 0.01 0.02 0.22 0.42 0.03 0.07
UMAP 0.61 0.73 0.23 0.15 0.52 0.74 0.74 0.82 0.33 0.36
t-SNE 0.71 0.77 0.17 0.17 0.41 0.70 0.73 0.80 0.21 0.27
PHATE 0.55 0.66 0.20 0.09 0.65 0.72 0.66 0.73 0.12 0.22
TriMap 0.59 0.72 0.2 0.14 0.62 0.77 0.74 0.81 0.15 0.2
PaCMAP 0.62 0.75 0.24 0.15 0.55 0.76 0.75 0.81 0.21 0.26
HeatGeo 0.6 0.69 0.15 0.10 N/A N/A 0.65 0.74 0.17 0.20
MS-IMAP 0.67 0.8 0.23 0.15 0.70 0.76 0.74 0.81 0.33 0.36
Table 1: Clustering results comparison using ARI and AMI on the Census, Zilionis, AWA, and Cora datasets. Note HeatGeo had code execution issues on the Zilionis dataset.
@@ 6 Feature Importance using Explicit Correspondence with Embedding Features
The correspondence established between the features and the embedding space offers a valuable tool for assessing feature importance. We explore two alternative methods for measuring feature importance, leveraging the constructed embedding space, where each dimension can be traced back to its corresponding feature in the original space. Specifically, we test these methods in an unsupervised or self-supervised setting where the subset of features are selected without using label information.
6.1 Assessing Embedding Feature Importance using the Laplacian Score
The first method utilizes the Laplacian Score (LS) [19], which assesses the importance of each feature by examining its correlation with the eigenvalues of the graph Laplacian. The significance of the variable corresponding to the $l$-th coordinate, represented by the feature $f_{l}\in\mathbb{R}^{N}$, is inferred from the relevance of its corresponding coordinate $l$ in the embedding space, $\tilde{\psi}_{\mathbf{x}_{l}}$. The Laplacian Score is calculated with respect to the embedding features $\tilde{\psi}_{\mathbf{x}}$ using the graph Laplacian $\mathbf{L}$ and the degree matrix $\mathbf{D}$. To compute the Laplacian Score, we first subtract the mean and then calculate it as follows: $L_{s}(\tilde{\psi}_{\mathbf{x}})_{l}=\frac{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{L}(\tilde{\psi}_{\mathbf{x}})_{l}}{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{D}(\tilde{\psi}_{\mathbf{x}})_{l}}$
Smaller scores indicate that the embedding feature $(\tilde{\psi}_{\mathbf{x}})_{l}$ has a greater projection onto the subspace of eigenvectors corresponding to the smallest eigenvalues, signifying higher importance concerning the global graph structure. Therefore, the feature importance of $(\tilde{\psi}_{\mathbf{x}})_{l}$ can be directly interpreted as the importance of the corresponding original feature.
@@ 6.2 Assessing Importance of Embedding Features using Mutual Information
In the second method, we propose a self-supervised approach that uses Mutual Information (MI) to estimate the importance of each dimension in the embedding space, which can then be explicitly linked to the original features. First, we apply $k$-means to cluster the MS-IMAP embedding space and use the resulting clusters as pseudo-labels. We then compute MI to quantify the shared information between features and these pseudo-labels, similar to how MI would be computed with ground truth labels. This step allows MI to measure how much each feature contributes to the cluster assignments, highlighting those most informative in distinguishing clusters. The MI-Based Feature Importance identify features that are informative about cluster assignments. Thus, masking these features would especially degrade the clustering performance. We summarize the MI-based unsupervised feature selection in Algorithm 2. In comparison to the MI based feature importance method, using the Laplacian score [19] we have that the embedding may have combined original features in a way that makes individual Laplacian Scores less indicative of feature importance. The Laplacian score prioritize neighborhood preservation, which might not directly translate to cluster discrimination efficacy. On the other hand, while the proposed MI based feature importance identifies features that are directly informative for distinguishing clusters, one of its possible limitations may be relaying on the quality of pseudo-labels, which may not align with true clusters.
  Input: MS-IMAP embeddings $\psi_{\mathbf{x}}$, number of clusters $k$
1  Output: MI scores for each feature
@@
[…]

### v5

@@ 3.1 Step 1: Feature Representation to encode multi-scale structure
3D-Tensor Encoding: The 3D-Tensor Encoding method concatenates the multi-scale representation for all features at a fixed scale, generating a matrix representation $\psi_{\mathbf{x}}(s_{j},:,:)\in\mathbb{R}^{D}\times\mathbb{R}^{N}$ for each scale $s_{j}$, encoding SGW coefficients across features and nodes. After concatenating all $f_{i}$ at a fixed scale $s_{j}$, we have
@@
where we designate the concatenation using $\left|\right|$, with $\mathbf{c}(\psi_{{f}_{i}}(s_{k},:,:),\psi_{{f}_{j}}(s_{k},:,:))=\psi_{{f}_{i}}(s_{k},:,:)\left|\right|\psi_{{f}_{j}}(s_{k},:,:)$. The optimization can be performed separately for each scale $s_{j}$, or jointly for all scales $s_{j}$ using the 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$ . We detail the proposed feature representation encoding methods in the pseudo code algorithm 1.
@@
Output: Initial Node Embeddings ${\psi}_{\mathbf{x}}$ .
@@
1Step 4: Initial Embedding construction:
@@
Step 5: Concatenate $\psi_{{f}_{r}}(s_{j},:)$ for r=1,,Dr=1,...,D in a matrix form $\psi_{\mathbf{x}}(s_{j},:,:)$ for each fixed scale $s_{j}$, representing all scales using 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
@@
3D-Tensor Based Optimization for graph Embedding Encoding the representation in a 3D tensor offers two key advantages. First, is the computational efficiency: the update rule at each iteration, associated with a pair of nodes, is performed jointly and simultaneously for all scales. This avoids the need for independent optimization per scale, which would significantly increase computational complexity. The second is optimization alignment: jointly optimizing across scales ensures alignment in the optimization process for all scales with respect to a fixed pair of nodes (positive or negative sample). This approach is more effective than optimizing scales independently, as it maintains consistency and improves overall performance. While SGW captures global graph structure, each embedding update remains feature-specific, preserving the one-to-one correspondence between input features and embedding dimensions, which is essential for interpretability. The encoded manifold representation is a 3D tensor $\psi_{\mathbf{x}}\in\mathbb{R}^{K}\times\mathbb{R}^{D}\times\mathbb{R}^{N}$.
An SGD update rule at iteration $t$, applied to the 3D tensor, is:
@@
[…]
We apply the optimization based on SGD with respect to each scale $s_{k}$, where the gradient of the loss is approximated by:
@@
Feature updates: Our implementation employs an optimization process that operates jointly across all scales using the 3D tensor, as described in the SGD update rule for iteration $t$ in Eq.(2). The optimization process uses SGD with positive and negative sampling for all scales $s_{k}$ given a pair of positive or negative pair of nodes $i,j$. At each iteration $t$, associated with either a positive sample (an edge) or a negative sample (a pair of unconnected nodes), the update rule from Eq.(2) is applied. The derivative of the loss function, as detailed in Eq.(4), is computed for all scales $(s_{1},...s_{K})$, with the sign in Eq.(2) depending on whether the sample is positive or negative. Once the optimization concludes, the final embedding is derived from the 3D tensor by summing over the scales, as specified in Eq.(6). Recall that the term $\psi_{\mathbf{x}}(s_{k},:,:)\in\mathbb{R}^{D\times N}$ is a slice of the 3D SGW tensor at scale $s_{k}$, representing spectral graph wavelet coefficients for all features across all nodes $N$. The specific vector $\psi_{\mathbf{x}}(s_{k},j,:)\in\mathbb{R}^{N}$ corresponds to the SGW coefficients at scale $s_{k}$ for input feature $j$ across all nodes. Note that in Eq.(2), the $j$-th embedding coordinate—aligned with input feature jj—is updated via SGD using two components: (1) the current embedding value and (2) a structural term involving $v_{ij}^{\psi_{\mathbf{x}}(s_{k},:,:)}$, which integrates the graph structure and multiscale information from the SGW tensor. At each iteration, updates for positive/negative node pairs are applied jointly across all scales, but each scale’s contribution is weighted by its energy concentration—ensuring that more informative scales have greater influence. We compute the final embedding by summing up all the optimized embeddings with respect to each scale $s_{k}$
@@
The dimensionality of the final embedding is $\tilde{\psi}_{\mathbf{x}}\in\mathbb{R}^{D}\times\mathbb{R}^{N}$, maintaining the same dimensionality as the original input, and thus there exists a one-to-one correspondence between the coordinates of the input features $\mathbf{x}$ and $\tilde{\psi}_{\mathbf{x}}$. This correspondence ensures that each dimension of the feature space is directly mapped to a single dimension in the embedding space $\tilde{\psi}_{\mathbf{x}}$. Our optimization process involves using SGD with positive and negative sampling, similar to recent graph embeddings methods ([33, 25]). Positive edge samples are associated with attraction (first term on the right hand side), while negative samples (second term on the right hand side) refer to a pair of nodes that are not connected in the graph, which create repulsion among dissimilar points.
Remark 1: While we adopt the SGW formulation by Hammond et al. ([18]), our framework is compatible with other spectral graph wavelet and multiscale representations.
Remark 2: While the manifold hypothesis suggests the data lie near a lower-dimensional manifold, this does not preclude learning meaningful structure in the ambient space. Our method incorporates high-frequency graph information via spectral graph wavelets, enhancing stability and local structure capture, without requiring expensive eigenspectrum computations as in Laplacian Eigenmaps or UMAP.
@@ 4 Feature Importance using Explicit Correspondence with Embedding Features
Feature importance is one of the simplest yet most fundamental tools for model interpretability. It associates input features with model outputs, highlighting which features most strongly influence the learned representations. We show that feature subsets selected from the MS-IMAP embedding lead to superior clustering performance compared to those selected from the input space using the same feature importance methods. Unlike competing approaches, MS-IMAP retains semantic alignment and supports direct correspondence between embedding dimensions and original input features. This makes MS-IMAP an interpretable alternative to black-box models. We empirically demonstrate that applying feature importance techniques to the MS-IMAP embedding yields more informative feature subsets for clustering than applying the same techniques directly to the raw input. This provides evidence of the semantic consistency preserved by our method—often absent in traditional nonlinear embeddings. We explore two alternative approaches for estimating feature importance, leveraging the explicit correspondence in the embedding space where each dimension maps back to a specific input feature. These methods are evaluated in an unsupervised or self-supervised setting, where feature selection is performed without access to label information.
4.1 Assessing Embedding Feature Importance using the Laplacian Score:
The first method utilizes the Laplacian Score (LS) ([19]), which assesses the importance of each feature by examining its correlation with the eigenvalues of the graph Laplacian. The significance of the variable corresponding to the $l$-th coordinate, represented by the feature $f_{l}\in\mathbb{R}^{N}$, is inferred from the relevance of its corresponding coordinate $l$ in the embedding space, $\tilde{\psi}_{\mathbf{x}_{l}}$. Computing the Laplacian score with respect to the embedding features $\tilde{\psi}_{\mathbf{x}}$. The Laplacian score is calculated with respect to embedding features $\tilde{\psi}_{\mathbf{x}}$ using the Laplacian graph $\mathbf{L}$ and the degree matrix $\mathbf{D}$. To compute the Laplacian Score, we first subtract the mean and then calculate it as follows:
$L_{s}(\tilde{\psi}_{\mathbf{x}})_{l}=\frac{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{L}(\tilde{\psi}_{\mathbf{x}})_{l}}{(\tilde{\psi}_{\mathbf{x}})_{l}^{T}\mathbf{D}(\tilde{\psi}_{\mathbf{x}})_{l}}$
Smaller scores indicate that the embedding feature $(\tilde{\psi}_{\mathbf{x}})_{l}$ has a greater projection onto the subspace of eigenvectors corresponding to the smallest eigenvalues, signifying higher importance concerning the global graph structure. Therefore, the feature importance of $(\tilde{\psi}_{\mathbf{x}})_{l}$ can be directly interpreted as the importance of the corresponding original feature. Although the graph is typically constructed using local neighborhoods (e.g., via k-nearest neighbors), the Laplacian operator reflects global smoothness over the entire data manifold.
@@ 4.2 Assessing Importance of Embedding Features using Mutual Information
We also propose a self-supervised approach that uses Mutual Information (MI) to estimate the importance of each dimension in the embedding space, which can then be explicitly linked to the original features. First, we apply $k$-means to cluster the MS-IMAP embedding space and use the resulting clusters as pseudo-labels. We then compute MI to quantify the shared information between features and these pseudo-labels, similar to how MI would be computed with ground truth labels. This step allows MI to measure how much each feature contributes to cluster assignments, highlighting those that are the most informative to distinguish clusters. The MI-Based Feature Importance identify features that are informative about cluster assignments. Thus, masking these features would especially degrade the clustering performance. We summarize the MI-based unsupervised feature selection in Algorithm 2.
4.3 Comparison: Laplacian Score (LS) vs. Feature Importance Based-Mutual Information
Input: MS-IMAP embeddings $\tilde{\psi}_{\mathbf{x}}$, number of clusters $k$
1Output: MI scores for each feature
@@
2      $\text{clusters}=\text{k-means}(\tilde{\psi}_{\mathbf{x}},k)$